Sean wants to change 40 liters of a 60% solution of liquid hydrogen by draining off some liters and replacing them with an equal amount of a 20% solution of liquid hydrogen. If he wants to get a final mix of 52% of liquid hydrogen, how many liters must he drain off?
Answer by robertb(4012)

40*0.60 - x*0.60 + x*0.20 = 40*0.52
==> 24 - 0.40x = 20.8
==> -0.40x = -3.2
==> x = 8 liters must be drained off.